# sDPO: Don’t Use Your Data All at Once

### v1

@@ 3.5 Target Model Initialization in sDPO
The target model in each step of sDPO is also initialized with $M_{t-1}$, the aligned model from the last step. This ensures that the final model in sDPO has undergone training with the same amount of data as the final model in DPO. On the other hand, one concern of such design choice is that it may become increasingly difficult to stabilize the training of the target model as the steps progress, since it has already undergone training with a decreasing learning rate schedule in the preceding steps. Thus, another option is to use the initial SFT base model $S$ as the target model for all steps of sDPO.
However, as shown in Fig. 3, initializing the target model as $S$ results in a much bigger initial loss than that of $M_{t-1}$, which could lead to an unstable training. The main reason is that DPO training is usually done where the reference and target models are the same. In contrast, initializing the target model as $S$ creates a differential in the reference and target models, which may be amplified as the steps progress. Thus, for stable training, initializing the target model as $M_{t-1}$ was chosen for sDPO.
@@ 4 Conclusion
We propose sDPO where we use the preference data in a stepwise way instead of all at once. We show that applying sDPO results in more performant models than DPO in terms of H4 score. We also empirically exhibit that sDPO results in more aligned reference models by comparing mean $\gamma_{\pi_{ref}}$.
@@
While we have demonstrated the effectiveness of employing different datasets in distinct stages of sDPO, identifying an optimal strategy for segmenting more intricate DPO data collections remains an area for further exploration. This task is particularly challenging due to the complexities within these datasets. Our approach, while promising, necessitates a more deeper understanding of dataset characteristics and their impact on the performance of sDPO.
Furthermore, our experiments predominantly utilized SOLAR 10.7B models, driven by the state-of-the-art performance at the time of experimentation along with its unique 10.7 billion parameter size. The unique size of SOLAR 10.7B models made the usage of sDPO more necessary as there are far fewer open source LLMs that can be adopted as reference models.
@@
Evaluating the efficacy of LLMs is an evolving challenge in the field. In our study, we primarily employed tasks from the Huggingface Open LLM Leaderboard as benchmarks for evaluation. While this provided comparative results, future research could benefit from incorporating a wider array of tasks and benchmarks. These could include tasks that judge actual human or strong AI preference alignment. Such additional evaluation would not only enhance the validity of our findings but also contribute to the broader discourse on LLM assessment methodologies.
@@
Appendix A Related Work
A.1 Large Language Models
Recent research has highlighted a "scaling law" in the field of context-based language models Kaplan et al. (2020); Hernandez et al. (2021); Anil et al. (2023), showing a proportional relationship between the size of the model plus the training data and the resulting performance improvements. Consequently, this has led to the advent of Large Language Models (LLMs). In contrast to earlier models, LLMs can perform in-context learning, which includes abilities such as zero-shot learning Radford et al. (2019) and few-shot learning Brown et al. (2020), allowing them to adapt and perform tasks without the need for weight adjustments. These emergent abilities of LLMs, absent in their smaller counterparts, signal a significant evolution in language model capabilities Wei et al. (2022).
A.2 Alignment Tuning
To address these issues, there have been proposals for supervised fine-tuning methodologies such as Rank Responses to align Human Feedback (RRHF) Yuan et al. (2023), Reward rAnked FineTuning (RAFT) Dong et al. (2023), and Direct Preference Optimization (DPO) Rafailov et al. (2023). These methods circumvent the intricacies inherent in reinforcement learning and have been shown to yield empirical results on par with RLHF. Notably, the DPO technique straightforwardly encourages the LLM to favor positive responses and discourage negative ones. DPO has been observed to yield performant learning outcomes, in spite of its uncomplicated training procedure.
Concurrent to our work, Yuan et al. (2024) have developed an iterative framework for generating new preference datasets and performing DPO training on the resulting datasets. They empirically demonstrated the superiority of their iterative framework in terms of AlpacaEval 2.0. In contrast, our work is complementary to the above in the sense that we focus on utilizing the current preference data and does not undergo new data generation. Thus, our method can also be applied to Yuan et al. (2024) by changing the DPO training part to using sDPO instead. We leave the above combination as an interesting future work. Additionally, the evaluation used in Yuan et al. (2024) is also different to ours as we utilize tasks from Open LLM Leaderboard whereas Yuan et al. (2024) uses AlpacaEval 2.0.

### v2

@@ 4.5 Target Model Initialization in sDPO
One option for target model initialization in sDPO is to use $S$, the initial SFT base model, for all steps. However, such initialization results in the final model trained with sDPO seeing less data than using DPO instead. Further, the target model and the reference model become more and more different as the steps progress, which is a deviation from the original DPO setup and risks losing the theoretical benefits of DPO.
To concretely investigate such potential issues, we visualize the loss curves for initializing the target model as $S$ in Figure 3. We observe that the initial loss value is much higher when compared to initializing the target model as $M_{t-1}$, i.e., the same as the reference model and adhering to the DPO convention. As using $M_{t-1}$ the target model means that each step of sDPO is using the same setup as DPO, the loss curves are much more stable and desirable. Thus, for stable training, initializing the target model as $M_{t-1}$ was chosen for sDPO.
4.6 Easy-to-Hard Data Partitioning
The effectiveness of the easy-to-hard data partitioning used in sDPO is demonstrated in Table 3. Note that we use OpenOrca as $D_{1}$ and Ultrafeedback Cleaned as $D_{2}$. As ‘SOLAR 10.7B + SFT + sDPO’, which uses the easy-to-hard partitioning, is more performant than‘SOLAR 10.7B + SFT + sDPO Rand.’, which uses random partitioning, the proposed easy-to-hard data partitioning is more effective for sDPO.
4.7 Additional Results on Generation Tasks
Model EQ Bench MT Bench
SOLAR 10.7B + SFT + sDPO 68.83 7.43
SOLAR 10.7B + SFT + DPO 61.02 7.35
SOLAR 10.7B + SFT 60.48 7.14
Table 4: Additional results on EQ Bench Paech (2023) and MT Bench Zheng et al. (2023), both of which are generation tasks that highly correlate with the Chatbot Arena ELO Zheng et al. (2023); Chiang et al. (2024). The best scores for both benchmarks are shown in bold.
In Table 4, we also report results for EQ Bench Paech (2023) and MT Bench Zheng et al. (2023) for the SFT base model and the models obtained by applying DPO and sDPO on the SFT base model.
For EQ Bench, we use the version without the revision prompt. We note that the EQ Bench requires the models to generate an answer that can be parsed with a pre-defined template for evaluation, which could be said to measure distinct capabilities of LLMs from the log-probability benchmarks shown in Table 2. While applying DPO only mildly improves the performance from the SFT base model, applying sDPO improves the performance significantly by over $+8\%$, indicating the effectivenss in which sDPO improves the generation capabilities compared to DPO.
As for MT Bench, we note that using sDPO achieves the best score of 7.43 amongst the compared models. Notably, applying sDPO to the SFT base model improves the MT Bench score by a non-trivial margin of $+0.29$. Applying DPO to the SFT base model also improves the MT Bench score, but not by more than that of applying sDPO.
@@ 5 Conclusion
We propose sDPO, an extension of DPO for aligning LLMs. Unlike traditional DPO, sDPO employs a stepwise approach, using subsets of preference data sequentially. This method leverages the aligned model from the previous step as the reference for the current step, ensuring progressively better alignment. Our experiments demonstrate that sDPO significantly outperforms conventional DPO in terms of both log-probability benchmarks such as ARC, HellaSWAG, MMLU, and TruthfulQA, as well as generation benchmarks such as EQ Bench and MT Bench. Additionally, sDPO enhances model alignment, as indicated by higher mean $\gamma_{\pi_{ref}}$ values, showing improved alignment with human preferences. The stepwise nature of sDPO simplifies the training process and aligns with curriculum learning principles, facilitating a structured optimization path. By using existing preference datasets more effectively, sDPO results in higher performance and better-aligned language models. This approach has the potential to transform alignment tuning, offering a robust framework for future research in LLMs.
@@
While we have demonstrated the effectiveness of employing easy-to-hard data partitioning of different datasets in distinct stages of sDPO, identifying a more performant strategy for segmenting more intricate preference data collections remains an area for further exploration.
Furthermore, our experiments predominantly utilized SOLAR 10.7B models, driven by the state-of-the-art performance at the time of experimentation along with its relatively 10.7 billion parameter size. Although as SOLAR 10.7B models are also based on the Llama-2 architecture with our results likely to transfer to other similar decoder only transformer models, more experiments using other models would be beneficial.
@@
Evaluating the efficacy of LLMs is an evolving challenge in the field. In our study, we primarily employed tasks from the Huggingface Open LLM Leaderboard as benchmarks for evaluation along with EQ Bench and MT Bench. While this provided comparative results, future research could benefit from incorporating a wider array of tasks and benchmarks. These could include tasks that judge actual human or strong AI preference alignment. Such additional evaluation would not only enhance the validity of our findings but also contribute to the broader discourse on LLM assessment methodologies.